OpenStudy (anonymous):

The system of equations is inconsistent. What are the missing values? 5x + ___ y=2 ___x + 3y = 8

OpenStudy (tkhunny):

If you do the SAME thing and get DIFFERENT answers, that is INCONSISTENT.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

inconsistent means there will not be a solution to the system of two lines. In other words, the two lines will not intersect, and are parallel.

OpenStudy (tkhunny):

Are "8" and "2" the SAME?

OpenStudy (anonymous):

An easy way to determine if two lines are parallel from the given form above... Ax + By = E Cx + Dy = F If the constants are a multiple of one another than the lines are parallel. If you can multiply, Ax+By by some number , and get Cx+Dy, the lines will be parallel.

OpenStudy (anonymous):

So if you let the first equation be... 5x + 1y = 2 you can see that the second equation will be 3 times the first, The first will have 1y, and the second has 3y. So you have to multiply the 5x by 3

OpenStudy (anonymous):

3*(5x + 1y) ---> 15x + 3y

OpenStudy (anonymous):

The first blank is 3 the second blank is 5 therefore the same things cannot add up to 2 and 8 at the same time X) hope this helped

OpenStudy (anonymous):

hmm makes sense

OpenStudy (anonymous):

it did thankyou this is one of the subjects i am bad it

OpenStudy (anonymous):

It's alright hope you understand it better now

OpenStudy (anonymous):

i do :)

OpenStudy (anonymous):

As long as the coefficients on X and Y in the second line are a multiple of the coefficients of X and Y in the second line. Using 3 and 5, means that they used a multiple of 1. keep that in mind ;)

OpenStudy (tkhunny):

\(5x + 3y = \;\;2\) \(\underbrace{5x + 3y} = \underbrace{8}\) \(SAME\;\;\;Different\) Inconsistent.